• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name septum_auto

    Log entry time 15:53:38 on May 2, 2012

    Entry number 381497

    keyword=Wrong septum current, please check!

    Left septum set current 100A is not correct, should be 361.847A! please check!
    Left septum read back current 96.8011A is not correct, should be 100A! please check!
    Right septum set current 100A is not correct, should be 474.699A! please check!




    septum alarm using halog_post [halog_post 1.1]

    A copy of this log entry has been emailed to: jixie@jlab.org